OpenStudy (anonymous):

The maximum and minimum value of f(x) = ax^2+bx+c is found when?

OpenStudy (anonymous):

a.) x=-b/2a b.) x=b^2-4ac c.) x=0 d.) the quadrartic formula e.) all the above

OpenStudy (anonymous):

Would it be A?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

ohh okay thank you but how would it be that?

hartnn (hartnn):

the vertex of ax^2+bx+c=0 is at x = -b/2a are you allowed to use calculus if you want to prove it?

OpenStudy (anonymous):

oh yeah that makes sense and just for future reference. how can I use calculus to prove it?

OpenStudy (welshfella):

You find the derivative of f(x) and equate it to 0 Then solve for x

hartnn (hartnn):

the maximum or minimum of f(x) occurs at f'(x) = 0 where f'(x) is the derivative of f(x) \((ax^2+bx+c)' = 0 \\ 2ax+b = 0 \\ x = \dfrac{-b}{2a}\)

OpenStudy (anonymous):

Ohh okay thank you!